OpenStudy (anonymous):

You can use displacement to find the volume of an irregular object, such as a stone. Suppose the tank shown is filled with water to a depth of 8 inches. A stone is placed in the tank so that it is completely covered, causing the water level to rise by 2 inches. The volume of the stone is ____ in3. (Enter only the number.)

OpenStudy (anonymous):

Want these hands?

OpenStudy (jacobciezki):

yea

OpenStudy (anonymous):

Where you is cuz ill mess you up call me a shredder

OpenStudy (anonymous):

yo sebastian yhu funna lose

OpenStudy (anonymous):

@jacobciezki youngthug07 likes yhu

OpenStudy (anonymous):

really ill kiss her hand

OpenStudy (jacobciezki):

what

OpenStudy (anonymous):

@Swvillarreal2

OpenStudy (anonymous):

@GingaTheNinja<3

OpenStudy (anonymous):

@domebotnos

OpenStudy (jacobciezki):

i pm you

OpenStudy (domebotnos):

this makes no sense... the stone could be 6 ?????// i have no idea

OpenStudy (anonymous):

That's all that is given!! I am so confusedD: @domebotnos

OpenStudy (domebotnos):

i can tell... so am i @ganeshie8

OpenStudy (jacobciezki):

why you block me

OpenStudy (domebotnos):

i didnt block u

OpenStudy (anonymous):

Ok I get this question

OpenStudy (jacobciezki):

her

OpenStudy (anonymous):

You can completely disregard the 8 inches of water. That is just to throw you off

OpenStudy (domebotnos):

hmm

OpenStudy (anonymous):

What you want to focus on is the 2 inches. we know the container is 12 inches by 24 inches

OpenStudy (anonymous):

The water raises by 2 inches

OpenStudy (anonymous):

Volume is h * l * w This is 2*12*24

OpenStudy (anonymous):

So 576 is your inches cubed

OpenStudy (jacobciezki):

yep it is he is right

OpenStudy (anonymous):

you are a life saver!! @Fuzure could you possibly help me with the second part?

OpenStudy (anonymous):

sure

OpenStudy (jacobciezki):

yea

OpenStudy (anonymous):

If the stone were a perfect sphere, what would be its diameter to the nearest hundredth of an inch?

OpenStudy (jacobciezki):

576 in³ is correct sphere volume = 576 = (4/3)πr³ = πd³/6 d³ = 6*576/π d ≈ 10.32" -------------

OpenStudy (anonymous):

ok so we know the volume is 576 cubic inches right? Volume of a sphere is 4/3 * pi * r^3

OpenStudy (anonymous):

Yea, jacobcieki is correct

OpenStudy (jacobciezki):

yes

OpenStudy (anonymous):

jacobciezki*

OpenStudy (jacobciezki):

yes

OpenStudy (anonymous):

Thank you guys! And @jacobciezki I think i blocked you on accident sorry!!

OpenStudy (anonymous):

Did I fix it?

OpenStudy (jacobciezki):

ok

OpenStudy (jacobciezki):

no

OpenStudy (jacobciezki):

you know how